# Application of AI to formal methods — an analysis of current trends

### v1

@@ 4.1 Employed search query
Finding a search query that would return a manageable amount of relevant contributions proved difficult. The main problem was harnessing the bandwidth of terms used in the AI and FM domains. Hereby, the difficulty lies in the often ambiguous use of terms like “formal method”, which can be used in the sense of formal methods as outlined in Section 2.2 or as a term to describe an only somewhat formal approach to a problem. The situation is even more difficult for AI as AI and ML are often used interchangeably. In contrast, essential publications may only use the specific name of an applied technique without naming the fields of AI or ML.
Another issue was the cross-pollution of our results with studies from the related research field of verification for AI and ML systems, which is essentially the other direction in which FM is applied to AI. This field is also a highly relevant topic and shares the majority of potential search key phrases, further inflating our results tally.
@@
produce multiple thousands of results which we deemed impossible to process in a reasonable amount of time and which contained many studies outside of our intended scope. From probing some results as well as personal experience, we also knew that many relevant publications tend not to use these high-level terms, but use nomenclature specific to their sub-community instead. For instance, research regarding the formal B method tends to simply use “B method” instead of “formal method” as keyword. On the other hand, we noticed how relevant publications seemed also to prefer to use the concrete names of the applied AI algorithms rather than stating they used AI in a superficial manner. Therefore, we were concerned that a search query that is too abstract would not be able to penetrate the relevant fields well enough.
As a result, we settled on a more complex query containing the precise terminology of specific algorithms. For this, we distinguish between two sets of terms, the AI-terms (Fig. 2(b)) and the FM-terms (Fig. 2(a)). Both correspond to a disjunction of a set of selected terms from the respective field. The terms themselves were selected based on previous experiences in the field following PICO Kitchenham et al. [10].
The final search query (Fig. 2(c)) now makes use of both term aggregates to produce a corpus of primary studies in a strict and targeted manner. That is, we only looked for studies that make use of at least one AI term and at least one FM term in their titles, in their abstracts, and in their keywords. The reasoning behind this is to get a good initial penetration of all key fields while relevant but missed entries are found in the later search stages via snowballing.
@@
Figure 2: Construction of the used search query. The query consists of a disjunction of search terms for AI and FM, respectively. The unified query (2(c)) enforces that at least one keyword from each subfield is present in the title, the abstract, and the keywords of a respective publication. An asterisk (*) indicates a wildcard character that can match with any sub-word. For instance, “solv*” matches with solver but also with solving.
@@ 4.2 Database search and processing
The search was conducted in the last quarter of 2023. We used four meta-search engines [^1] Springer’s engine was not used because by the time of the search it performed extremely poor in dealing with large search queries, large result sets and filtering. as suggested by Petersen et al. [8]: IEEE [^2] https://ieeexplore.ieee.org/Xplore/home.jsp, Scopus  [^3] https://www.scopus.com/search/form.uri?display=basic#basic, ACM [^4] https://dl.acm.org/ and Web of Science (WoS) [^5] https://www.webofknowledge.com/. We used the Guide to Computing Literature for ACM. Furthermore, as the ACM search engine limits the number of wild cards, we split the search terms into subsets, performed the sub-searches, and merged them into the required superset. For WoS, results seemed to differ depending on which institution has access. Therefore, we decided to take the more extensive result set. As far as the search engines allowed, we applied the EC, such as the area of publishing and the publication language. The result size after this step was 1492 entries. This number and all later numbers are without duplicates.
@@ 4.3 Inclusion and exclusion criteria
For the 1492 resulting contributions, we made two observations. First, the amount was too large for a snowballing procedure to be feasible. Second, while briefly looking at the corpus, we discovered many contributions that should not have been selected, i.e., contributions that included the term “specification” but meant it in a purely requirement engineering-minded context. Therefore, we decided to apply our IC and EC prematurely to the corpus as far as they applied, i.e., to title and abstract. For this, we read the titles of our studies, and if we could not decide whether the contributions should be included, we also conducted the abstracts, and very rarely, the contributions itself was skimmed.
To achieve this for such a large corpus, the first and second authors passed over all entries individually and decided whether they should be included or excluded. In agreement cases, the respective studies were kept or discarded, respectively. For cases of disagreement, both authors discussed each instance together to find an agreement. The disagreement could not be settled this way in two cases, and the third author was consulted as a tiebreaker. This procedure brought the number of primary studies down to 89 but took a significant amount of time.
@@ 1. EC 1:
    The contribution appears to be not reviewed in any form.
@@ 3. EC 3:
    The contribution not posed in computer science.
@@ 6. EC 6:
    (During initial skim) The contribution focuses only on common or general constrain-solving techniques (without specializing in SMT/SAT)
@@ 10. EC 10:
For the EC, some additional criteria were added in later steps in response to contributions encountered that were of low value for this study. For instance, EC 6 was added because many contributions focused on constraint solving, thus being closer to mathematics. EC 7 was added in response to a large influx of SAT contributions that mentioned learning but focused on remembering already-seen clauses. EC 8 is a rule that was introduced to not have secondary studies in the result set but to use those studies to ensure that even with a restrictive query, we cover as much of the field as possible. EC 10 was added in the fifth snowballing round, where we got many contributions that synthesized automatons, but no further analysis was mentioned. EC 9 was added as we found 2Sat as a problem not relevant for this study, as solutions are achievable in polynomial time already.
@@ 4.4 Snowballing
For the 89 entries, we conducted an extensive snowballing procedure [46] where we conducted both forward and backward snowballing. Five iterations were needed until we eventually reached a closure. This extensive snowballing approach was meant to complement the strict search query and to uncover relevant yet initially missed studies or even subfields of research. Therefore, snowballing served as a means to alleviate these threats to validity.
For the backward snowballing, we consulted the references of a given contribution and selected promising titles or publications that were explicitly highlighted in the respective related work sections. If a publication appeared promising but the contribution was not clear from its title alone, the abstract was consulted as well. For the forward snowballing, we used Google Scholar and applied the same title and abstract procedure. This brought the number of contributions up to 540. After concluding the snowballing, we applied the IC and EC a second time, reducing the amount of contributions to 457.
@@ 5 Results
In the following section, we will discuss the results. For this, we will answer the individual research questions and aim to make a cross-analysis between individual research questions, thus extracting as much information as possible from the existing data set. We refer to Section 6 for a discussion of the observed results.
@@ 5.1 RQ 1: Demographics of the research area
The whole corpus of 457 primary studies we found was published in the years 1972–2023. Figure 3 displays the histogram of publications per year (see also Appendix A) and highlights that AI applications on FM appear to mirror the trend of general AI application mentioned in Section 1. That is, we can observe a steadily growing interest in the field, especially since around 2005.
Focusing on the last five years only (recall Section 4.5), there still appears to be growing interest overall, while not each year surpasses the respective previous one. For 2019, we have 32, for 2020, we have 48, for 2021, we have 39, for 2022, we have 51, and for 2023, we have 21 contributions. Noteworthy are two observations: First, for the last five years alone, there has been no strict upward trend in the number of publications, We can observe a decline from 2020 to 2021. Second, there were few publications in 2023. We explain this because we searched in the last quarter of 2023 when not all possible studies were published. Given that we do not yet have complete data for 2023, it is impossible to forecast whether the overall trend will continue to grow or if the current peak from 2022 marks a global maximum.
@@
Figure 4 (see also Appendix A) displays the publications over the years divided by publication venues. The majority of studies were published as conference papers while journal papers seem to be on the decline again since their peak in 2018.
Within our five-year observation range, 122 contributions were conference papers, followed by 39 journal articles 29 workshop or short papers, and only one (1) book chapter with no full books.
Figure 4: Distribution of publication venues
5.1.3 RQ 1.3: What are the provided main contributions by the primary studies?
Following the definitions of contribution types from Section 2.4, we see the division of the studies into these contribution types in Fig. 5(a) (see also Appendix A). We can see that the majority of contributions were methodologies (118), i.e., practical application of AI to FM, followed by studies with multiple types of contributions (29). Third place was tool contributions (14), followed by improvements for tools (10), followed by frameworks (6), training data (5), case studies (2), benchmarks (2), approaches (2), and finally one (1) idea contribution. Studies with multiple contribution types most commonly provide a methodology (20/29), training data (14/29), or a tool (11/29). See Fig. 5(b) for a complete breakdown.
@@
We can gain some additional insights by investigating the individual results in context to each other. Figure 6 (see also Appendix A) shows how the publication venues and the contribution types correlate to each other. We can see that for any contribution type, a heavy focus lies on conference papers. Interestingly, we can see that case studies are only conducted within the scope of journal papers and book chapters. Furthermore, workshops and short papers are primarily used to introduce new methodologies, which seem unexpected, as usually, due to limited space, authors aim to restrict themselves to outlining an approach or an idea.
@@ 5.2 RQ 2: Content type of contributions
In Fig. 7(a) (see also Appendix B), we can see that the majority of contributions use NN (70) followed by RL (32). 27 contributions use multiple techniques, while 16 use NLP methods and 14 use EA. A total of 8 studies presented custom algorithms. The rest of the contributions are divided between utilizing decision trees (7), clustering (3), Bayesian inference (3), KNN (2), random forests (2), data mining approaches (2), automaton learning (2), and SVM (1).
Studies that employed multiple algorithms mostly did so in a contrasting manner, i.e., conducted training on numerous different models to see which one performed best for their respective application. A detailed view of the category of multiple algorithms is given in Fig. 7(b) (see also Appendix B). Here, we can see that if multiple techniques are used, they mainly rely on NN (13/31), random forest (12/31), and some variant of tree learners (12/31). Interestingly, SVM (6/31) and KNN (5/31) applications are more predominant in a setting with multiple algorithms compared to being the sole focus of a study. We explain this by KNN or SVM being simple enough algorithms for authors to train them in a small amount of time. Further, they do need less fine-tuning than, say, NNs, lending themselves as baseline approaches for comparisons.
@@
Figure 7: Overview of AI contributions. Underrepresented AI types are aggregated as Other with a number in parenthesis indicating how many sub-techniques were included.
@@
Figure 8 (see also Appendix B) shows the intersection of RQs 2.1 and 1.3. The wide focus on methodologies was already assessed with Fig. 5(a). In Fig. 8, we can see that the amount of NN contributions remained focused on methodologies (37/189) or as part of a broader application of techniques (13/189) or a tool (8/189).
@@
Figure 9 (see also Appendix B) show the intersection of RQs 2.1 and 1.1. Here, we can see that the amount of NN contributions remained steady over the years, similar to the amount of RL contributions. Noteworthy is that there is no notable decline in any technique over the observed period. The alternating nature of the overall amount of contributions was already found back in Section 5.1.
@@
Figure 10(a) (see also Appendix B) gives an overview of the FM techniques found in the contributions. TP is leading by a large quantity with 85 entries. SAT is next with 45 entries. Synthesis approaches come in third place with 19 entries, followed by model checking (18) and SMT solving (13). The remaining entries fall into algorithm selection, program analysis, and termination analysis.
In Fig. 10(b) (see also Appendix B), we can see a more detailed investigation of the topic of theorem proving. Here, we can see that premise, axiom, and clause selection received the most attention (27/85), followed by proof search (21/85) and proof synthesis (18/85). After that, a significant gap exists in a row of smaller topics.
In Fig. 10(c), we see the division of SAT approaches into their application areas. Overall, 37/45 studies focused on typical SAT solving, 2 on QBFs, and 6 on 3SAT. Out of 45 articles, 9 focused on solving, 9 on finding MaxSAT, 7 on solver selection, 6 on feature extraction, 5 aimed to predict solubility and four contributions aimed at analyzing the performance of solvers. A small group of contributions aimed to do multiple things generated SAT problems and developed branching heuristics aimed at parameter selection or dependency analysis.
Figure 10(d) (see also Appendix B) takes a closer look at the topic of synthesis. 8/19 contributions aimed to synthesize a model or specification, 4/19 aimed to learn a loop invariant, 3/19 aimed to repair models, and 3/19 targeted general invariant learning. One contribution targets the generation of annotations for the verification of JML code.
@@
(b) Quantity of targeted theorem proving topics
@@
Figure 10: Quantity of FM approaches with a detailed view on the most relevant contributions types
@@
A closer inspection is warranted as TP makes up the most significant portion of contributions. For the studies concerned with TP, we distinguished between higher-order logic (47/85) and first-order logic (38/85). We further classified them between their automation levels: fully automatic TP (74/85), interactive TP (10/85), and contributions employing both (1/85). This granularity gives rise to additional observations about the problem structure. Figure 11(a) (see also Appendix B) shows that the majority of contributions (26/85) were made in the area of premise selection for automatic theorem proving, followed by proof search (21/85) and proof synthesis (18/85).
Figure 11(b) (see also Appendix B) shows that of the 38 first-order entries, there is a tight focus on premise selection (20/38), proof search (8/38) and proof synthesis (7/38). Although not as drastic, a similar spread can be observed for the higher-order contributions. Premise selection has 7/47 entries, proof search 13/47, and proof synthesis 11/47 entries. 4/47 entries focused on tactics prediction.
@@
(b) Automatization grade against the type of TP subcategory
Figure 11: Theorem proving in details
@@
Figure 12(a) (see also Appendix B) shows the different types of FM over the years. We can see no particular concentration on one year, similar to Fig. 9, However, there may be some trends. SAT had two solid years in 2020 and 2022, while TP had a slow decline; contributions to model checking and synthesis increased. Nonetheless, given the small time frame of five years, Any actual trends might be invisible to us.
In Fig. 12(b) (see also Appendix B), we showcase the distribution of FM types over contribution types. While all FM techniques found AI applications in the last five years, we now see a fragmented distribution of specific contribution types per discipline. In general, for TP, the different types of available contributions are the broadest, while for model checking, we are restricted to methodologies and one tool. The notable absence of tool improvements outside of TP might indicate developments in the respective subfields. There might be no developments that deem an extension of tools necessary, or no relevant tools can be extended. However, the underlying reasons for this observation are out of the scope of this mapping study. For the synthesis of formal models, we see a slightly better situation regarding a broader landscape of contribution types. Overall, however, the absence of case studies and benchmarks deserves attention and might highlight that the field is still in the early stages of development.
@@
(b) Distribution of FM techniques over contribution types. Studies with multiple contribution types are considered for each distinct contribution type individually.
@@
In Fig. 13, we can see several overlays of FM applications with FM techniques in different granularity. Figure 13(a) (see also Appendix B) is the most abstract. Here, we can see that the bulk of the theorem-proving techniques apply NNs (38), RL (16), and NLP (12). SAT solvers mainly utilize NNs (24). Model checking frequently uses EA (6) and NN (5). The area of synthesis is mainly divided between NNs (5), RL (4), and NLP (6) as well.
Taking a closer look at TP in Fig. 13(b) (see also Appendix B), we can see that NN is primarily used for the premise, axiom, and clause selection (18) proof search (7) and proof synthesis (10) RL techniques are often used for proof search as well (8). With three exceptions for heuristic selection, premise selection, or proof search, EA is absent from the area of theorem proving.
Focusing on SAT in Fig. 13(c) (see also Appendix B), we can see that NN is dominant in most subcategories, except for portfolio selection and branching heuristics. EA has higher relevance for MaxSat.
Finally, taking a closer look at the synthesis topic in Fig. 13(d), we can see that NLP is relevant for the model generation (see also Appendix B) and the generation of verification annotations. NN and RL are mainly used for (loop) invariant learning. For the repair of formal models, the contributions consider multiple techniques.
@@
(b) Theorem proving techniques with applied AI techniques
@@
Figure 13: Quantity of FM approaches and comparison with time frame and contribution type. Studies that applied multiple AI techniques types are considered for each distinct AI technique individually. Underrepresented techniques are aggregated as Other with a number in parenthesis indicating how many sub-techniques were included.

### v2

@@ 4.1 Employed search query
Finding a search query that would return a manageable amount of relevant contributions proved difficult. The main problem was harnessing the bandwidth of terms used in the AI and FM domains. Hereby, the difficulty lies in the often ambiguous use of terms like “formal method”, which can be used in the sense of formal methods as outlined in Section 2.2 or as a term to describe an only somewhat formal approach to a problem. The situation is even more difficult for AI, as AI and ML are often used interchangeably. In contrast, essential publications may only use the specific name of an applied technique without naming the fields of AI or ML.
Another issue was the cross-pollution of our results with studies from the related research field of verification for AI and ML systems. This is essentially the other direction in which FM is applied to AI. This field is also a highly relevant topic and shares the majority of potential search key phrases, further inflating our results tally.
@@
produce multiple thousands of results which we deemed impossible to process in a reasonable amount of time and which contained many studies outside of our intended scope. From probing some results as well as personal experience, we also knew that many relevant publications tend not to use these high-level terms, but use nomenclature specific to their sub-community. For instance, research regarding the formal B method tends to simply use “B method” instead of “formal method” as keyword. On the other hand, we noticed how relevant publications also seemed to prefer to use the concrete names of the applied AI algorithms, rather than stating they used AI in a superficial manner. Therefore, we were concerned that a search query that was too abstract would not be able to penetrate the relevant fields well enough.
As a result, we settled on a more complex query containing the precise terminology of specific algorithms. For this, we distinguish between two sets of terms, the AI-terms (Fig. 2(b)) and the FM-terms (Fig. 2(a)). Both correspond to a disjunction of a set of selected terms from the respective field. The terms themselves were selected based on previous experiences in the field following the PICO (Population Intervention Comparision Outcome) approach as suggested by Kitchenham et al. [10].
Here Population may refer to a specific field of FM or AI, such as formal model(ing) or deep learn(ing). Intervention may refer to specific techniques or methodologies of FM or AI, such as SAT or SVM. Comparision was not applicable as we focused on a direct connection, while Outcomes shall contain at least strongly correlated AI and FM terms.
The final search query (Fig. 2(c)) now uses both term aggregates to produce a corpus of primary studies in a strict and targeted manner. We only looked for studies that use at least one AI term and at least one FM term in their titles, abstracts, and keywords. The reasoning is to get a good initial penetration of all key fields, while relevant but missed entries are found in the later search stages via snowballing.
@@
Figure 2: Construction of the used search query. The query consists of a disjunction of search terms for AI and FM, respectively. The unified query (2(c)) enforces that at least one keyword from each subfield is present in the title, the abstract, and the keywords of a respective publication. An asterisk (*) indicates a wildcard character that can match any sub-word. For instance, “solv*” matches with solver but also with solving.
@@ 4.2 Database search and processing
The search was conducted in the last quarter of 2023. We used four meta-search engines [^1] Springer’s engine was not used because by the time of the search it performed extremely poor in dealing with large search queries, large result sets and filtering. as suggested by Petersen et al. [8]: IEEE [^2] https://ieeexplore.ieee.org/Xplore/home.jsp, Scopus  [^3] https://www.scopus.com/search/form.uri?display=basic#basic, ACM [^4] https://dl.acm.org/ and Web of Science (WoS) [^5] https://www.webofknowledge.com/. We used the Guide to Computing Literature for ACM. Furthermore, as the ACM search engine limits the number of wildcards, we split the search terms into subsets, performed the subsearches, and merged them into the required superset. For WoS, results seemed to differ depending on which institution had access. Therefore, we decided to take the more extensive result set. As far as the search engines allowed, we applied the EC, such as the area of publishing and the publication language. The result size after this step was 1492 entries. This number and all later numbers are without duplicates.
@@ 4.3 Inclusion and exclusion criteria
For the 1492 resulting contributions, we made two observations. First, the amount was too large for a snowballing procedure to be feasible. Second, while briefly looking at the corpus, we discovered many contributions that should not have been selected, i.e., contributions that included the term “specification” but meant it in a purely requirement engineering-minded context. Therefore, we decided to apply our IC and EC prematurely to the corpus as far as they applied, i.e., to the title and abstract. For this, we read the titles of our studies. If we could not decide whether the contributions should be included, we also conducted an abstract review, and very rarely, the contributions themselves were skimmed.
To achieve this for such a large corpus, the first and second authors passed over all entries individually and decided whether they should be included or excluded. In agreement cases, the respective studies were kept or discarded. For disagreement, both authors discussed each instance together to reach an agreement. The disagreement could not be settled this way in two cases, and the third author was consulted as a tiebreaker. This procedure reduced the number of primary studies to 89, but took time.
@@ 1. EC 1:
    The contribution appears not to have been reviewed in any form.
@@ 3. EC 3:
    The contribution is not posed in computer science.
@@ 6. EC 6:
    (During initial skim) The contribution focuses only on common or general constraint-solving techniques (without specializing in SMT/SAT)
@@ 10. EC 10:
For the EC, some additional criteria were added in later steps in response to contributions encountered that were of low value for this study. For instance, EC 6 was added because many contributions focused on constraint solving, thus being closer to mathematics. EC 7 was added in response to a large influx of SAT contributions that mentioned learning but focused on remembering already-seen clauses. EC 8 is a rule that was introduced to not have secondary studies in the result set but to use those studies to ensure that even with a restrictive query, we cover as much of the field as possible. EC 10 was added in the fifth snowballing round, where we got many contributions that synthesized automatons, but no further analysis was mentioned. While an automaton may represent a formal model, analysis of larger automatons is very human-unfriendly compared to mathematical formulas or formal models written in a modeling language. Therefore, we expected the authors to subject the generated automaton to some validation, e.g., model checking or SAT/SMT solving, to enable reasoning about the validity of the generated automaton. EC 9 was added as we found 2Sat as a problem irrelevant for this study, as solutions are already achievable in polynomial time.
@@ 4.4 Snowballing
For the 89 entries, we conducted an extensive snowballing procedure [45] where we conducted both forward and backward snowballing. Five iterations were needed until we eventually reached closure. This extensive snowballing approach was meant to complement the strict search query and to uncover relevant yet initially missed studies or even subfields of research. Therefore, snowballing served as a means to alleviate these threats to validity.
For the backward snowballing, we consulted the references of a given contribution and selected promising titles or publications that were explicitly highlighted in the respective related work sections. If a publication appeared promising, but the contribution was unclear from its title alone, the abstract was also consulted. For the forward snowballing, we used Google Scholar and applied the same title and abstract procedure. This brought the number of contributions up to 540. After concluding the snowballing, we applied the IC and EC a second time, reducing the number of contributions to 457.
@@ 5 Results
In the following section, we will discuss the results. For this, we will answer the individual research questions and aim to make a cross-analysis between individual research questions, thus extracting as much information as possible from the existing dataset. We refer to Section 6 for a discussion of the observed results. An overview of the results matched to their respective application domain in FM can be found in Appendix A.
@@ 5.1 RQ 1: Demographics of the research area
The whole corpus of 457 primary studies we found was published in the years 1972–2023. Figure 3 displays the histogram of publications per year and highlights that AI applications on FM appear to mirror the trend of general AI applications mentioned in Section 1. That is, we can observe a steadily growing interest in the field, especially since around 2005.
Focusing on the last five years only (recall Section 4.5), there still appears to be growing interest overall, while not each year surpasses the respective previous one. We have 32 contributions for 2019, 48 for 2020, 39 for 2021, 51 for 2022, and 21 for 2023. Noteworthy are two observations: First, for the last five years alone, there has been no strict upward trend in the number of publications, We can observe a decline from 2020 to 2021. Second, there were few publications in 2023. We explain this because we searched in the last quarter of 2023, when not all possible studies were published. Given that we do not yet have complete data for 2023, it is impossible to forecast whether the overall trend will continue to grow or if the current peak from 2022 marks a global maximum.
@@
Figure 4 displays the publications over the years, divided by publication venues. The majority of studies were published as conference papers while journal papers seem to be on the decline again since their peak in 2018. Within our five-year observation range, 124 contributions were conference papers, followed by 37 journal articles 29 workshop or short papers, and only one (1) book chapter with no full books.
Figure 5 shows the publication venues sorted by the main targeted area. For this, we skimmed through the titles of the conferences and sorted them into six different areas. We can see the majority of publications (94) happened in AI-focused venues, 49 contributions were published in software engineering venues, 40 in FM venues, 19 in venues concerned with automated reasoning specifically, 8 in mathematics-focused venues, and 20 in other, less topical venues. Note that some venues have more than one area of focus, e.g., we counted publications from the conference on artificial intelligence and theorem proving (AITP) for both AI and FM.
Figure 4: Distribution of publication venues by submission type
Figure 5: Distribution of publication venues by area
5.1.3 RQ 1.3: What are the main contributions provided by the primary studies?
Following the definitions of contribution types from Section 2.4, we see the division of the studies into these contribution types in Fig. 6(a). We can see that the majority of contributions were methodologies (118), i.e., practical application of AI to FM, followed by studies with multiple types of contributions (29). Third place was tool contributions (14), followed by improvements for tools (10), followed by frameworks (6), training data (5), case studies (2), benchmarks (2), approaches (2), and finally, one (1) idea contribution. Studies with multiple contribution types most commonly provide a methodology (20/29), training data (14/29), or a tool (11/29). See Fig. 6(b) for a complete breakdown.
@@
We can gain some additional insights by investigating the individual results in context with each other. Figure 7 shows how the publication venues and the contribution types correlate. We can see that for any contribution type, a heavy focus lies on conference papers. Interestingly, we can see that case studies are only conducted within the scope of journal papers and book chapters. Furthermore, workshops and short papers are primarily used to introduce new methodologies, which seem unexpected, as usually, due to limited space, authors aim to restrict themselves to outlining an approach or an idea.
@@ 5.2 RQ 2: Content type of contributions
In Fig. 8(a), we can see that the majority of contributions use NN (70), followed by RL (32). 27 contributions use multiple techniques, while 16 use NLP methods and 14 use EA. A total of 8 studies presented custom algorithms. The rest of the contributions are divided between utilizing decision trees (7), clustering (3), Bayesian inference (3), KNN (2), random forests (2), data mining approaches (2), automaton learning (2), and SVM (1).
Studies that employed multiple algorithms mostly did so in a contrasting manner, i.e., they trained on numerous models to see which one performed best for their respective applications. A detailed view of the category of multiple algorithms is given in Fig. 8(b). Here, we can see that if multiple techniques are used, they mainly rely on NN (13/31), random forest (12/31), and some variant of tree learners (12/31). Interestingly, SVM (6/31) and KNN (5/31) applications are more predominant in a setting with multiple algorithms compared to being the sole focus of a study. We explain this by KNN or SVM, which are simple algorithms that authors can train quickly. Further, they need less fine-tuning than NNs, making them a baseline comparison approach.
@@
Figure 8: Overview of AI contributions. Underrepresented AI types are aggregated as Other with a number in parentheses indicating how many sub-techniques were included.
@@
Figure 9 shows the intersection of RQs 2.1 and 1.3. The wide focus on methodologies was already assessed with Fig. 6(a). In Fig. 9, we can see that the amount of NN contributions remained focused on methodologies (37/189) or as part of a broader application of techniques (13/189) or a tool (8/189).
@@
Figure 10 show the intersection of RQs 2.1 and 1.1. Here, we can see that the amount of NN contributions remained steady over the years, similar to the amount of RL contributions. Noteworthy is that there is no notable decline in any technique over the observed period. The alternating nature of the overall amount of contributions was already found back in Section 5.1.
@@
Figure 11(a) gives an overview of the FM techniques in the contributions. TP is leading by a large margin with 85 entries. SAT is next with 45 entries. Synthesis approaches come in third place with 19 entries, followed by model checking (18) and SMT solving (13). The remaining entries fall into algorithm selection, program analysis, and termination analysis.
In Fig. 11(b), we can see a more detailed investigation of the topic of theorem proving. Here, we can see that premise, axiom, and clause selection received the most attention (27/85), followed by proof search (21/85) and proof synthesis (18/85). After that, a significant gap exists in a row of smaller topics.
In Fig. 11(c), we see the division of SAT approaches into their application areas. Overall, 37/45 studies focused on typical SAT solving, 2 on QBFs, and 6 on 3SAT. Out of 45 articles, 9 focused on solving, 8 on finding MaxSAT, 7 on solver selection, 6 on instance selection, 5 aimed to predict solvability, and 4 contributions aimed at analyzing the performance of solvers. A small group of contributions aimed to do multiple things, generated SAT problems, and developed branching heuristics aimed at parameter selection or dependency analysis.
Figure 11(d) takes a closer look at the topic of synthesis. 8/19 contributions aimed to synthesize a model or specification, 4/19 aimed to learn a loop invariant, 3/19 aimed to repair models, and 3/19 targeted general invariant learning. One contribution targets the generation of annotations for the verification of JML code.
@@
(b) Quantity of targeted theorem-proving topics
@@
Figure 11: Quantity of FM approaches with a detailed view of the most relevant contribution types
@@
A closer inspection is warranted, as TP makes up the most significant portion of contributions. For the studies concerned with TP, we distinguished between higher-order logic (47/85) and first-order logic (38/85). We further classified them by their automation levels: fully automatic TP (74/85), interactive TP (10/85), and contributions employing both (1/85). This granularity gives rise to additional observations about the problem structure. Figure 12(a) shows that the majority of contributions (26/85) were made in the area of premise selection for automatic theorem proving, followed by proof search (21/85) and proof synthesis (18/85).
Figure 12(b) shows that of the 38 first-order entries, there is a tight focus on premise selection (20/38), proof search (8/38), and proof synthesis (7/38). Although not as drastic, a similar spread can be observed for the higher-order contributions. Premise selection has 7/47 entries, proof search 13/47, and proof synthesis 11/47 entries. 4/47 entries focused on tactics prediction.
@@
(b) Automation grade against the type of TP subcategory
Figure 12: Theorem proving in detail
@@
Figure 13(a) shows the different types of FM over the years. We can see no particular concentration on one year, similar to Fig. 10, However, there may be some trends. SAT had two solid years in 2020 and 2022, while TP had a slow decline; contributions to model checking and synthesis increased. Nonetheless, given the small time frame of five years, any actual trends might be invisible to us.
In Fig. 13(b), we showcase the distribution of FM types over contribution types. While all FM techniques have found AI applications in the last five years, we now see a fragmented distribution of specific contribution types per discipline. In general, for TP, the different types of available contributions are the broadest, while for model checking, we are restricted to methodologies and one tool. The notable absence of tool improvements outside of TP might indicate developments in the respective subfields. There might be no developments that deem an extension of tools necessary, or no relevant tools can be extended. However, the underlying reasons for this observation are out of the scope of this mapping study. For the synthesis of formal models, we see a slightly better situation regarding a broader landscape of contribution types. Overall, however, the absence of case studies and benchmarks deserves attention and might highlight that the field is still in the early stages of development.
@@
(b) Distribution of FM techniques over contribution types. Studies with multiple contribution types are considered for each distinct contribution type.
@@
In Fig. 14, we see several overlays of FM applications with FM techniques at different granularities. Figure 14(a) is the most abstract. Here, we can see that the bulk of the theorem-proving techniques apply NNs (38), RL (16), and NLP (12). SAT solvers mainly utilize NNs (24). Model checking frequently uses EA (6) and NN (5). The area of synthesis is mainly divided between NNs (5), RL (4), and NLP (6) as well.
Taking a closer look at TP in Fig. 14(b) , we can see that NN is primarily used for the premise, axiom, and clause selection (17) proof search (7) and proof synthesis (10). RL techniques are often used for proof search as well (8). With three exceptions for heuristic selection, premise selection, or proof search, EA is absent from the area of theorem proving.
Focusing on SAT in Fig. 14(c), we can see that NN is dominant in most subcategories, except for portfolio selection and branching heuristics. EA has higher relevance for MaxSat.
Finally, taking a closer look at the synthesis topic in Fig. 14(d), we can see that NLP is relevant for the model generation and the generation of verification annotations. NN and RL are mainly used for (loop) invariant learning. For the repair of formal models, the contributions consider multiple techniques.
@@
(b) Theorem-proving techniques with applied AI techniques
@@
Figure 14: The quantity of FM approaches and comparison with a time frame and contribution type. Studies that applied multiple AI techniques are considered individually for each distinct AI technique. Underrepresented techniques are aggregated as Other with a number in parentheses indicating how many sub-techniques were included.